OpenStudy (jjuden):

A right cone has a base diameter that is 6 cm and a slant height that is 27 cm. What is the surface area of the cone?

OpenStudy (jjuden):

\[A.81\pi cm^2 B.84\pi cm^2 C.90\pi cm^2 D. 162\pi cm^2\]